# Equivariant category of wedges

### v4

@@ 3. Equivariant Category of Wedges
First, we show that for any $G$-connected, $G$-normal space with a $G$-non-degenerate basepoint, there exists a “good” open $G$-categorical cover. This claim is the equivariant version of [5, Lemma 1.25, pg. 13].
@@
Without loss of generality, we can assume that $x_{0}\in U_{j}$, for $j=1,\ldots,k$, for some $11\leq k\leq n-$ 1, and $x_{0}\notin U_{j}$, for $j=k+1,\ldots,n$. Define a new $G$-contracting open invariant neighborhood of $x_{0}$ by
@@
$[(U_{j}\cap(X-\overline{M}))\cup M]\cup\mathcal{N}\supset U_{j}.$
@@
$[(U_{j}\cap(X-\overline{M}))\cup M]\cup\mathcal{N}\supset W_{j},$
@@
For $G$-spaces $X$ and $Y$, we will assume that $G$ acts diagonally on $X\times Y$, meaning that for all $(x,y)\in X\times Y$ and for all $g\in G$, we have $g(x,y):=(gx,gy)$. If $x_{0}\in X^{G}$ and $y_{0}\in Y^{G}$, the wedge sum $X\vee Y=X\times\{y_{0}\}\cup\{x_{0}\}\times Y\subset X\times X$ is an invariant subset of the $G$-space $X\times Y$. Indeed, for any $x\in X$, $y\in Y$, and any $g\in G$, we have $g(x,y_{0})=(gx,gy_{0})=(gx,y_{0})\in X\vee Y$ and $g(x_{0},y)=(gx_{0},gy)=(x_{0},gy)\in X\vee Y$. Hence, $X\vee Y$ is a $G$-space. Furthermore, note that for any $x\in X$ and $y\in Y$, the orbits in $X\vee Y$ are $O(x,y_{0})=O(x)\times\{y_{0}\}$ and $O(x_{0},y)=\{x_{0}\}\times O(y)$. In addition, observe that the wedge $X\vee Y$ coincides with the space obtained from the disjoint union $X\sqcup Y$ by identifying the basepoints $x_{0}$ and $y_{0}$.
We have that the inequality $\max\{\text{cat}_{G}(X),\text{cat}_{G}(Y)\}\leq\text{cat}_{G}(X\vee Y)$ always holds.
@@
Then, we state one of the main purposes of this paper. This claim is the equivariant version of [5, Proposition 1.27, pg. 14]. In the next section, we will use this theorem to compute the equivariant and invariant topological complexities for $G$-connected $G$-CW-complexes with $X^{G}\neq\varnothing$.
@@
The statement that $X/A$ is $G$-homotopy equivalent to $X\vee SA$ follows by analogy to [10, Example 0.14, p.14]. Hence, by Proposition 2.5, we obtain that $\text{cat}_{G}(X/A)=\text{cat}_{G}(X\vee SA)$. From Theorem 3.4, we conclude that $\text{cat}_{G}(X/A)=\text{cat}_{G}(X)$, since $\text{cat}_{G}(X)\geq 2$ (using the fact that $X$ is not $G$-contractible) and $\text{cat}_{G}(SA)\leq 2$ (see Example 2.3(3)). ∎
@@
For $X,Y$ $G$-spaces, we recall that $G$ acts diagonally on $X\times Y$. If $x_{0}\in X^{G}$ and $y_{0}\in Y^{G}$, the wedge $X\vee Y=X\times\{y_{0}\}\cup\{x_{0}\}\times Y\subset X\times X$ is an invariant subset of the $G$-space $X\times Y$, and the smash product
@@
is a $G$-space. We do not know the value of the equivariant category $\text{cat}_{G}(X\wedge Y)$ for any pointed $G$-spaces $X$ and $Y$. However, for $Y=S^{k}$ ($k\geq 0$) with the trivial action of $G$, we have the following statement.
Let $H_{q}(-;\mathbb{Z})$ denote the $q$th singular homology group. We have the following statement.
@@
Let $X$ be a pointed $G$-space such that $H_{q}(X;\mathbb{Z})\neq 0$, for some $q\geq 1$, and let $S^{k}$ be the $k$-dimensional sphere ($k\geq 0$) with the trivial action of $G$, such that $X\wedge S^{k}$ is $G$-connected (for $k\geq 1$). Then, we have
@@
For the case $k\geq 1$, it is also a technical procedure to check that there exists a $G$-equivariant homeomorphism between $X\wedge S^{k}$ and the iterated reduced suspension $\Sigma^{k}X$. Hence, by Proposition 2.5, we have $\text{cat}_{G}(X\wedge S^{k})=\text{cat}_{G}(\Sigma^{k}X)$. Note that $\text{cat}_{G}(\Sigma^{k}X)\leq 2$. Then, by Corollary 2.11, we have
@@
$\leq$ $2$
The first inequality follows from the fact that $H_{q}(X;\mathbb{Z})\neq 0$, for some $q\geq 1$ (using the Mayer–Vietoris sequence, cf. [16, Lemma 4.10, p 36]). ∎
@@ 4. Equivariant and Invariant Topological Complexities
Let $X$ be a $G$-space. Let $X^{[0,1]}$ denote the space of all continuous paths in $X$, equipped with the compact-open topology, and let $e_{2}^{X}:X^{[0,1]}\to X\times X$ be the fibration given by $e_{2}^{X}(\gamma)=\left(\gamma(0),\gamma(1)\right)$. Note that $X^{[0,1]}$ is a $G$-space via the pointwise action $(g\gamma)(t):=g\gamma(t)$, for any $g\in G$, $\gamma\in X^{[0,1]}$, and $t\in[0,1]$. The Cartesian product $X\times X$ is also a $G$-space via the diagonal action $g(x,y)=(gx,gy)$. Hence, $e_{2}^{X}:X^{[0,1]}\to X\times X$ is a $G$-map (in fact, it is a $G$-fibration, see [4, p. 2308]).
@@
Note that $G\times G$ acts on $X^{[0,1]}\times_{X/G}X^{[0,1]}$ via $(g,h)(\alpha,\gamma):=(g\alpha,h\gamma)$, and on $X\times X$ via $(g,h)(x,y)=(gx,hy)$. Hence, the map $p_{X}:X^{[0,1]}\times_{X/G}X^{[0,1]}\to X\times X$, given by $p_{X}(\alpha,\gamma):=(\alpha(0),\gamma(1))$, is a $G\times G$-map (indeed, it is a $G\times G$-fibration, see [13, Proposition 3.7, p. 110]).
@@
    [4, Definition 5.1, p. 2309] The equivariant topological complexity of $X$, denoted by $\text{TC}_{G}(X)$, is the smallest integer $k$ such that $X\times X$ can be covered by $G$-invariant open sets $U_{1},\ldots,U_{k}\subset X\times X$, each which admitting a $G$-map $s_{i}:U_{i}\to X^{[0,1]}$ such that $e_{2}^{X}\circ s_{i}=\text{incl}_{U_{i}}$. If no such integer $k$ exists, then we set $\text{TC}_{G}(X)=\infty$.
@@
    [13, p. 110] The invariant topological complexity of $X$, denoted by $\text{TC}^{G}(X)$, is the smallest integer $k$ such that $X\times X$ may be covered by $G\times G$-invariant open sets $U_{1},\ldots,U_{k}\subset X\times X$, each which admitting a $G\times G$-map $s_{i}:U_{i}\to X^{[0,1]}\times_{X/G}X^{[0,1]}$ such that $p_{X}\circ s_{i}=\text{incl}_{U_{i}}$. If no such integer $k$ exists, then we set $\text{TC}^{G}(X)=\infty$.
@@
    [4, Theorem 5.2, p. 2309], [13, Proposition 3.25, p. 115] If $X$ $G$-dominates $Y$, then $\text{TC}_{G}(X)\geq\text{TC}_{G}(Y)$ and $\text{TC}^{G}(X)\geq\text{TC}^{G}(Y)$. In particular, if $X\simeq_{G}Y$, then $\text{TC}_{G}(X)=\text{TC}_{G}(Y)$ and $\text{TC}^{G}(X)=\text{TC}^{G}(Y)$.
@@
    [4, Corollary 5.8, p. 2310], [13, Remark 3.24, p. 114] If $X$ is a $G$-connected $G$-CW-complex with $X^{G}\neq\varnothing$, then
@@
    [4, Corollary 5.4(1), p. 2310], [13, Corollary 3.26, p. 115] If $X$ is a $G$-space, then
@@
Theorem 3.7 shows that the equality $\text{cat}_{G}(X/A)=\text{cat}_{G}(X)$ holds under the condition that the inclusion map $A\hookrightarrow X$ is $G$-homotopic to a constant map $\overline{x_{0}}:A\to X$ for some $x_{0}\in X^{G}$. We can obtain similar formulas for the equivariant and invariant topological complexities under the condition that $A$ is $G$-contractible to a fixed point $x_{0}\in X^{G}$. For this purpose, we show the equivariant version of [10, Proposition 0.17, p. 15]. As a direct consequence, we compute the equivariant category, and the equivariant and invariant topological complexities of $X/A$.
@@
Now, we present the equivariant and invariant version of [7, Theorem 3.6, p. 4371].
@@
As another application of Theorem 3.4 together with Theorem 4.4, we obtain the following statement, which is the equivariant and invariant case of [17, Proposição 4.7.53, p. 177].
@@
We present the following example. For completeness, let us first recall that for $n\geq 2$, $S^{n}$ is the $n$-sphere with the group $\mathbb{Z}_{2}$ acting by reflection, given by multiplication by $-1$ in the last coordinate. We have $\left(S^{n}\right)^{\mathbb{Z}_{2}}=S^{n-1}$, and $\text{TC}_{\mathbb{Z}_{2}}(S^{n})=2\text{cat}_{\mathbb{Z}_{2}}(S^{n})-1=3$, as shown in [4, Example 5.9, p. 2311]. On the other hand,
$\text{TC}^{\mathbb{Z}_{2}}(S^{n})=\begin{cases}2,&\hbox{ if $n$ is even,}\\ 3,&\hbox{ if $n$ is odd,}\\ \end{cases}$
as shown in [13, Example 4.2, p. 115]. Also, given integers $n_{1},\ldots,n_{k}\geq 1$ with $k\geq 2$, we have $\text{TC}\left(S^{n_{1}}\vee\cdots\vee S^{n_{k}}\right)=3$, as shown in [17, Proposição 4.7.50, p. 176].
@@
$\text{TC}_{G}(X\wedge S^{k})=\text{TC}_{G}(\Sigma^{k}X)\quad\text{ and }\quad\text{TC}^{G}(X\wedge S^{k})=\text{TC}^{G}(\Sigma^{k}X).$
@@
This conjecture, in the nonequivariant case, was proved in [8, Theorem 6, p. 409] under more restrictive hypotheses.

### v5

@@ 3. Equivariant Category of Wedges
First, we show that for any $G$-connected, $G$-normal space with a $G$-non-degenerate basepoint, there exists a “good” open $G$-categorical cover. This claim is the equivariant version of [6, Lemma 1.25, pg. 13].
@@
Without loss of generality, we can assume that $x_{0}\in U_{j}$, for $j=1,\ldots,k$, for some $1\leq k\leq n-1$, and $x_{0}\notin U_{j}$, for $j=k+1,\ldots,n$. Define a new $G$-contracting open invariant neighborhood of $x_{0}$ by
@@
$(U_{j}\cap(X-\overline{M}))\cup M]\cup\mathcal{N}\supset U_{j}.$
@@
$(U_{j}\cap(X-\overline{M}))\cup M]\cup\mathcal{N}\supset W_{j},$
@@
For $G$-spaces $X$ and $Y$, we will assume that $G$ acts diagonally on $X\times Y$, meaning that for all $(x,y)\in X\times Y$ and for all $g\in G$, we have $g(x,y):=(gx,gy)$. If $x_{0}\in X^{G}$ and $y_{0}\in Y^{G}$, the wedge sum $X\vee Y=X\times\{y_{0}\}\cup\{x_{0}\}\times Y\subset X\times X$ is an invariant subset of the $G$-space $X\times Y$. Furthermore, note that for any $x\in X$ and $y\in Y$, the orbits in $X\vee Y$ are $O(x,y_{0})=O(x)\times\{y_{0}\}$ and $O(x_{0},y)=\{x_{0}\}\times O(y)$. In addition, observe that the wedge $X\vee Y$ coincides with the space obtained from the disjoint union $X\sqcup Y$ by identifying the basepoints $x_{0}$ and $y_{0}$.
@@
Then, we state one of the main purposes of this paper. This claim is the equivariant version of [6, Proposition 1.27, pg. 14]. In the next section, we will use this theorem to compute the equivariant and invariant topological complexities for $G$-connected $G$-CW-complexes with $X^{G}\neq\varnothing$.
@@
The statement that $X/A$ is $G$-homotopy equivalent to $X\vee SA$ follows by analogy to [11, Example 0.14, p.14]. Hence, by Proposition 2.5, we obtain that $\text{cat}_{G}(X/A)=\text{cat}_{G}(X\vee SA)$. From Theorem 3.4, we conclude that $\text{cat}_{G}(X/A)=\text{cat}_{G}(X)$, since $\text{cat}_{G}(X)\geq 2$ (using the fact that $X$ is not $G$-contractible) and $\text{cat}_{G}(SA)\leq 2$ (see Example 2.3(3)). ∎
@@
For the $G$-spaces $X$ and $Y$, we recall that $G$ acts diagonally on $X\times Y$. If $x_{0}\in X^{G}$ and $y_{0}\in Y^{G}$, the wedge $X\vee Y=X\times\{y_{0}\}\cup\{x_{0}\}\times Y\subset X\times X$ is an invariant subset of the $G$-space $X\times Y$, and the smash product
@@
is a $G$-space. The value of the equivariant category $\operatorname{cat}_{G}(X\wedge Y)$ is not known for arbitrary pointed $G$-spaces $X$ and $Y$. However, when $Y=S^{k}$ (for $k\geq 0$) equipped with the trivial $G$-action, we have the following result.
@@
Let $H_{q}(-;\mathbb{Z})$ denote the $q$th singular homology group. Let $X$ be a pointed $G$-space such that $H_{q}(X;\mathbb{Z})\neq 0$, for some $q\geq 1$, and let $S^{k}$ be the $k$-dimensional sphere ($k\geq 0$) equipped with the trivial $G$-action. Assume that $X\wedge S^{k}$ is $G$-connected for $k\geq 1$. Then:
@@
For the case $k\geq 1$, it is also a technical procedure to check that there exists a $G$-equivariant homeomorphism between $X\wedge S^{k}$ and the iterated reduced suspension $\Sigma^{k}X$. Hence, by Proposition 2.5, we have $\text{cat}_{G}(X\wedge S^{k})=\text{cat}_{G}(\Sigma^{k}X)$. Note that $\text{cat}_{G}(\Sigma^{k}X)\leq 2$. Then, we have
@@
$\leq$ $2.$
The first inequality follows from the fact that $H_{q}(X;\mathbb{Z})\neq 0$, for some $q\geq 1$ (using the Mayer–Vietoris sequence, cf. [18, Lemma 4.10, p 36]). ∎
@@ 4. Equivariant and Invariant Topological Complexities
Let $X$ be a $G$-space. Let $X^{[0,1]}$ denote the space of all continuous paths in $X$, equipped with the compact-open topology, and let $e_{2}^{X}:X^{[0,1]}\to X\times X$ be the fibration given by $e_{2}^{X}(\gamma)=\left(\gamma(0),\gamma(1)\right)$. Note that $X^{[0,1]}$ is a $G$-space via the pointwise action $(g\gamma)(t):=g\gamma(t)$, for any $g\in G$, $\gamma\in X^{[0,1]}$, and $t\in[0,1]$. The Cartesian product $X\times X$ is also a $G$-space via the diagonal action $g(x,y)=(gx,gy)$. Hence, $e_{2}^{X}:X^{[0,1]}\to X\times X$ is a $G$-map (in fact, it is a $G$-fibration, see [5, p. 2308]).
@@
Note that $G\times G$ acts on $X^{[0,1]}\times_{X/G}X^{[0,1]}$ via $(g,h)(\alpha,\gamma):=(g\alpha,h\gamma)$, and on $X\times X$ via $(g,h)(x,y)=(gx,hy)$. Hence, the map $p_{X}:X^{[0,1]}\times_{X/G}X^{[0,1]}\to X\times X$, given by $p_{X}(\alpha,\gamma):=(\alpha(0),\gamma(1))$, is a $G\times G$-map (indeed, it is a $G\times G$-fibration, see [14, Proposition 3.7, p. 110]).
@@
    [5, Definition 5.1, p. 2309] The equivariant topological complexity of $X$, denoted by $\text{TC}_{G}(X)$, is the smallest integer $k$ such that $X\times X$ can be covered by $G$-invariant open sets $U_{1},\ldots,U_{k}\subset X\times X$, each which admitting a $G$-map $s_{i}:U_{i}\to X^{[0,1]}$ such that $e_{2}^{X}\circ s_{i}=\text{incl}_{U_{i}}$. If no such integer $k$ exists, then we set $\text{TC}_{G}(X)=\infty$.
@@
    [14, p. 110] The invariant topological complexity of $X$, denoted by $\text{TC}^{G}(X)$, is the smallest integer $k$ such that $X\times X$ may be covered by $G\times G$-invariant open sets $U_{1},\ldots,U_{k}\subset X\times X$, each which admitting a $G\times G$-map $s_{i}:U_{i}\to X^{[0,1]}\times_{X/G}X^{[0,1]}$ such that $p_{X}\circ s_{i}=\text{incl}_{U_{i}}$. If no such integer $k$ exists, then we set $\text{TC}^{G}(X)=\infty$.
@@
    [5, Theorem 5.2, p. 2309], [14, Proposition 3.25, p. 115] If $X$ $G$-dominates $Y$, then $\text{TC}_{G}(X)\geq\text{TC}_{G}(Y)$ and $\text{TC}^{G}(X)\geq\text{TC}^{G}(Y)$. In particular, if $X\simeq_{G}Y$, then $\text{TC}_{G}(X)=\text{TC}_{G}(Y)$ and $\text{TC}^{G}(X)=\text{TC}^{G}(Y)$.
@@
    [5, Corollary 5.8, p. 2310], [14, Remark 3.24, p. 114] If $X$ is a $G$-connected $G$-CW-complex with $X^{G}\neq\varnothing$, then
@@
    [5, Corollary 5.4(1), p. 2310], [14, Corollary 3.26, p. 115] If $X$ is a $G$-space, then
@@
Theorem 3.7 shows that the equality $\text{cat}_{G}(X/A)=\text{cat}_{G}(X)$ holds under the condition that the inclusion map $A\hookrightarrow X$ is $G$-homotopic to a constant map $\overline{x_{0}}:A\to X$ for some $x_{0}\in X^{G}$. We can obtain similar formulas for the equivariant and invariant topological complexities under the condition that $A$ is $G$-contractible to a fixed point $x_{0}\in X^{G}$. For this purpose, we show the equivariant version of [11, Proposition 0.17, p. 15]. As a direct consequence, we compute the equivariant category, and the equivariant and invariant topological complexities of $X/A$.
@@
Now, we present the equivariant and invariant version of [8, Theorem 3.6, p. 4371].
@@
As another application of Theorem 3.4 together with Theorem 4.4, we obtain the following statement, which is the equivariant and invariant case of [19, Proposição 4.7.53, p. 177].
@@
We present the following example. For completeness, let us first recall that for $n\geq 2$, $S^{n}$ is the $n$-sphere with the group $\mathbb{Z}_{2}$ acting by reflection, given by multiplication by 1-1 in the last coordinate. We have $\left(S^{n}\right)^{\mathbb{Z}_{2}}=S^{n-1}$, and $\text{TC}_{\mathbb{Z}_{2}}(S^{n})=2\text{cat}_{\mathbb{Z}_{2}}(S^{n})-1=3$, as shown in [5, Example 5.9, p. 2311]. On the other hand,
[…]
as shown in [14, Example 4.2, p. 115]. Also, given integers $n_{1},\ldots,n_{k}\geq 1$ with $k\geq 2$, we have $\text{TC}\left(S^{n_{1}}\vee\cdots\vee S^{n_{k}}\right)=3$, as shown in [19, Proposição 4.7.50, p. 176].
@@
[…]
@@
This conjecture, in the nonequivariant case, was proved in [9, Theorem 6, p. 409] under more restrictive hypotheses.